OpenStudy (anonymous):

find the measure of the three angles of a triangle when the smallest angle is 20 less than the second angle and the third angle is 50 less than three times the second angle

OpenStudy (ja1):



OpenStudy (anonymous):

Solve following equations for a, b and c. a = b-20, c = 3b-50, a+b+c=180 Verify that a=30, b=50 and c=100 .

OpenStudy (anonymous):

@JA1 Thank you for the medal.

OpenStudy (ja1):

@robtobey well you did answer the question correctly :)